No relevant resource is found in the selected language.

This site uses cookies. By continuing to browse the site you are agreeing to our use of cookies. Read our privacy policy>Search

Reminder

To have a better experience, please upgrade your IE browser.

upgrade
Knowledge Base

When the OSN 1800 optical cable is broken, the logical fiber on the NMS does not turn red.

Publication Date:  2019-07-21  |   Views:  148  |   Downloads:  0  |   Author:  lWX689240  |   Document ID:  EKB1100022552

Contents

Issue Description

NMS version: U2000V2R17C60SPC200,OSN 1800 host version V1R6C00SPC300.The main control board is F3SCC, which is configured with the F1FIU01 board,The OSC_LOS alarm is reported on the OCH path,The receive-end OPU connected to the FIU01 also reports the MUT_LOS alarm But he FIU board is not report any LOS alarm,As a result, the logical fibers between NEs are not red.The FIU board does not suppress or mask the MUT_los alarm.

Handling Process

Local NE FIU board does not report MUT-LOS alert causes logical fiber not to turn red,The FIU board does not report the MUT_LOS alarm because the peer OSN 1800 NE uses the F1SCC control board and works with the F1FIU board.The two boards do not support fiber connection setting. Therefore, the MUT_LOS alarm is not supported.

Solution

The workaround is to skip the FIU board for logical fiber connections between NEs on the NMS.